# Improving the Quality of Web-mined Parallel Corpora of Low-Resource Languages using Debiasing Heuristics

### v2

@@ 5.3 Human Evaluation Results
As shown in Table 2, heuristic-based PDC had reduced the noise in the top-ranked samples consistently, irrespective of the language pair and the considered multiPLM. Some gains are quite significant - For example, the amount of correct pairs (C) improvement for CCMatrix-EnSi (LaBSE), CCAligned-EnTa (LaBSE) and CCAligned-EnSi (XLM-R) were all 82%. Before applying the heuristics, the average Error category (E) across language pairs and datasets is 13.50% for LASER3, 36.33% for XLM-R and 72.53% for LaBSE. CS, CCN and UN noise categories were noted to be contributing towards this noise percentage.
However, after heuristic filtration, the error drops drastically. For LASER, XLM-R and LaBSE on average, it is 2.17%, 2.50% and 1.0%, respectively. However, the untranslated (UN) and overlapping text (CCN) were still there. To eliminate these types of noise, we would need to employ an alignment model similar to Steingr´ımsson et al. (2023) or Minh-Cong et al. (2023).
In conclusion, the human evaluation results also indicate that the heuristic-based PDC approach is beneficial for multiPLM ranking in two key ways. First, it produces the top-ranked sentence pairs from multiPLM to be qualitatively comparable. Secondly, it removes the noisy parallel sentences that cause the disparity among the NMT systems trained using ranked corpora based on multiPLMs.
@@ 6 NMT Performance based on Training Data Size
We further investigate the effect of the training dataset size on the final NMT performance. We sample 50k, 100k, 150k, and 200k from each ranked corpus, and train NMT models. Results are in Figure 5 of the Appendix H.
For Si-Ta, moderate but consistent improvements were observed from 50K to 100K samples for both CCMatrix and CCAligned ranked corpora, irrespective of the multiPLMs. However, due to the lack of sufficient corpus size, the NMT scores for 150K and 200K were not observable. For En-Ta, NMT scores declined after 100K, while for En-Si, an improvement was observed until 200K, except for the CCMatrix XLM-R ranked sample. Hence, we suspect that noise may be more prominent beyond 100K for En-Ta. In conclusion, increasing corpus size can enhance NMT performance, but only when the underlying data is sufficiently clean and well-aligned.
7 Discussion
We show that applying simple rule-based heuristics to remove noisy sentences before using a multiPLM is better than simply ranking the web-mined corpora based on embeddings derived from multiPLMs as suggested by Ranathunga et al. (2024).
Based on our extensive experiments, dedup+punctNums+(n)gram+sLength+LIDThresh
+sentWRatio is our recommended heuristic combination, based on the justifications provided earlier. Although $n$ can be corpus dependent, we recommend starting with n=5 or 6. In contrast to the recent PDC work Steingr´ımsson (2023); Minh-Cong et al. (2023) that employs several deep learning model-based rigorous filtration pipelines, our technique is much simpler.
@@ 8 Conclusion
In this research, we empirically analysed the disparity between the NMT systems trained with the web-mined corpora ranked using embeddings derived from multiPLMs. With a human evaluation, we showed that this disparity is due to different types of noise creeping into the top-ranked portion of corpora when different multiPLMs are used. We made use of rule-based heuristics to remove this noise. After a systematic evaluation of heuristics, we were able to identify optimal heuristic combinations that resulted in higher NMT scores. Therefore, for anyone planning to use web-mined corpora, our recommendation is to first filter out noisy sentences using heuristics and then to do ranking on the embeddings derived from the multiPLM. Human evaluation indicated that even after applying heuristics, some noise remains. In future, we plan to apply techniques such as classification-based approaches to remove such noise, or to apply translation post-editing to improve sentence quality.
@@ 9.1 Limitations
We found that the LID was suboptimal for identifying Si or Ta languages. Therefore, such models would need to be optimised or better LID models would need to be used, such that they are effective in predicting the language label correctly. Due to the lack of funding, we were able to employ only two annotators for the task. However, the results would have been more convincing if the annotation could have been done by three annotators. Furthermore, due to the lack of human annotators, we had to limit this analysis to only three language pairs. We can extend this study to more low-resource language pairs provided that we can find suitable annotators.
@@ 9.2 Ethical Concerns
We use publicly available datasets. Due to the dataset size, we did not have the resources or funding to check all the curated parallel sentences manually for offensive content. Fernando et al. (2020) provided the human-crafted dataset to be used in this research. Further, we do not disclose or share any personal details of the annotators publicly or with any other party. We have offered the annotators the standard rate and have settled the payments in full. The only details we disclose are in the Appendix F.
@@ Acknowledgments
Will be provided upon acceptance.
Appendix A Improved Noise Taxonomy
The definitions of all the noise categories in our improved noise taxonomy are given in Table 9. Then, in Table 10, we show examples of parallel sentences which fall under these noise categories.
Error Code: Description
@@ Appendix A Improved Noise Taxonomy
CCN: Overlapping Content
@@
Table 9:  Ranathunga et al. (2024)’s error taxonomy with the CCN category that has been newly added by us
Table 10: Example parallel sentences from the En-Si, En-Ta and Si-Ta, identified during human evaluation.
Appendix B Selection of Languages and Datasets
Sinhala is an Indo-Aryan language spoken primarily in Sri Lanka by the Sinhalese majority. It exhibits complex morphological structures, including rich inflectional and derivational processes, but is classified as a low-resource language due to the scarcity of linguistic resources and tools.
Tamil, a Dravidian language with a rich literary history, is spoken by Tamil communities in Sri Lanka, India, and the global diaspora. Unlike Sinhala, Tamil benefits from a relatively larger digital presence, but it still faces challenges in NLP applications due to morphological complexity, agglutinative grammar, and resource limitations in certain domains.
We obtain the publicly released CCMatrix and CCAligned datasets from the OPUS website [^13] https://opus.nlpl.eu/. Both these datasets support the language pairs, En-Si, En-Ta and Si-Ta, which we consider in this research.
Appendix C Selection of multiPLMs
LASER3 Heffernan et al. (2022) (L=12, H=1024, A=4, 250M) [^15] No of Layers, Hidden Layer Dimensions and No of Attention Heads are defined by L, H and A respectively. is a multiPLM favourable for bitext mining and cross-lingual tasks. It improves over previous LASER2 versions by supporting more languages and enhancing alignment quality, but still faces challenges in low-resource settings.
XLM-R Conneau et al. (2020) (L=12, H=768, A=6, 278M parameters) is a transformer-based multiPLM trained on massive amounts of text using masked language modelling. It achieves strong cross-lingual performance but struggles with low-resource languages due to limited training data.
LaBSE Feng et al. (2022) (L=12, H=768, A=12, 471M) is a BERT-based model optimized for multilingual sentence embeddings and bitext retrieval. It provides high-quality cross-lingual representations and is favourable for cross-lingual tasks.
Appendix D Sentence Length Threshold
To determine the optimal threshold for sentence length filtering, we filter sentences less than 3, 4, and 5 words on the source (S) side, target (T) side, and both sides. Similar to other PDC experiments, we then rank the parallel sentence pairs by computing embeddings using three different multiPLMs, and train NMT models using the top 100k ranked data. We conduct these experiments using the CCAligned corpus, for En-Si and En-Ta language-pairs. The ChrF++ results are shown in Table 11. We observe consistently sentence length threshold of 5 yields the best results. Therefore, we select this threshold in all subsequent experiments.
@@ Appendix D Sentence Length Threshold
Appendix E NMT Experiments
The experiments were conducted on a NVIDIA Quadro RTX6000 GPU with 24GB VRAM. The hyperparameters used during training, along with the training parameters, are shown in Table 12. We conduct training on the NMT systems for 100 epochs with early stopping criteria and report the results using ChrF++. ChrF++ was chosen over the conventional multi-BLEU Papineni et al. (2002) and sacreBLEU (Post, 2018) because character-level evaluation is more suitable for the considered Sinhala and Tamil languages, which are morphologically rich in nature.
@@ Appendix E NMT Experiments
Appendix F Human Evaluation
Annotator Experience Qualification
Annotator 01 22 Diploma in Translation And Interpretation
Annotator 02 9 BSc.(Hons) in Information Technology
Annotator 03 5 MBBS
Annotator 04 4 BA (Hons) in Translation
Annotator 06 2.5 Diploma in Translation And Interpretation
Annotator 07 2.5 BSc Eng (Hons) Electrical & Electronics Engineering
Annotator 09 2 Bachelor of Industrial Information Technology
Resources Provided and Training: All the annotators have had prior experience with a similar task. However, for this annotation work, we provide them with the definitions of the noise categories (Table 9) along with example sentence pairs (Table 10) and the guideline in terms of a flowchart (Figure 3). First, we asked them to do a sample of 30 sentences as a training on the task, and review it. Then, the 1200 sentence pairs to be annotated were shared with each annotator via Google Sheets to be completed.
Compensation: They were paid for each sentence pair they annotated, which was the standard rate for such work.
Appendix G Results Analysis
We show the individual and the combined heuristic which produced the best NMT gains with respect to each multiPLM, dataset and the language pair in the Table 14. The Figure 4 shows the percentage of dedup+ngram experiments exceeding the highest dedup, with respect to the language-pair. Finally, we show the final dataset sizes after applying the heuristic along with the percentage reduction in Table 15.

### v3

@@ 5.3 Human Evaluation Results
As shown in Table 2, heuristic-based PDC had reduced the noise in the top-ranked samples consistently, irrespective of the language pair and the considered multiPLM. Some gains are quite significant — for example, the amount of correct pairs (C) improvement for CCMatrix-EnSi (LaBSE), CCMatrix-EnTa (LaBSE), and CCAligned-EnSi (LaBSE) were 82%, 82%, and 74% respectively. CS, CCN and UN noise categories are noted to be contributing towards this noise percentage. However, after heuristic filtration, the error drops drastically. On average, for LASER, XLM-R, and LaBSE, the final error percentages are 2.17%, 2.50% and 1.0%, respectively. The evaluation further reveals that the residual noise are from untranslated (UN) and overlapping text (CCN) classes. These findings reveal that rule-based heuristics, are not effective in eliminating those types of noise. As a result, we would need to employ an alignment model similar to Steingrímsson et al. (2023) or Minh-Cong et al. (2023) to remove such residual noise from the corpus.
In conclusion, the human evaluation results indicate that the heuristic-based PDC approach is beneficial for parallel sentence ranking in two key ways. First, it produces the top-ranked sentence pairs from multiPLM to be qualitatively comparable. Secondly, it removes the noisy parallel sentences that cause the disparity among the NMT systems trained using ranked corpora based on multiPLMs.
@@ 6 NMT Performance based on Training Data Size
We further investigate the effect of the training dataset size on the final NMT performance. We sampled the top 50k, 100k, 150k, and 200k from each of the ranked corpora after applying the heuristic-combination, and train NMT models. Results are in Figure 5 of the Appendix H.
For Si-Ta, moderate but consistent improvements are observed from 50K to 100K samples for both CCMatrix and CCAligned corpora, irrespective of the multiPLMs. However, after heuristic-based filtering, the resulting Si-Ta parallel corpus contained only about 100K sentence pairs (Table 15 in Appendix G). Therefore, NMT experiments with 150K and 200K sentence pairs could not be conducted due to insufficient data.
For En-Ta, NMT scores declined after 100K, while for En-Si, an improvement was observed until 200K, except for the CCMatrix ranked using XLM-R embeddings. Hence, we suspect that noise may be more prominent beyond 100K for En-Ta. In conclusion, increasing corpus size can enhance NMT performance, but only when the underlying data is sufficiently clean and well-aligned.
@@ 7 Conclusion
In this research, we empirically analysed the disparity between the NMT systems trained with the web-mined corpora ranked using embeddings derived from multiPLMs. With a human evaluation, we showed that this disparity is due to different types of noise creeping into the top-ranked portion of corpora when different multiPLMs are used.
We made use of rule-based heuristics to remove this noise. After a systematic evaluation of heuristics, we were able to identify optimal heuristic combinations that resulted in higher NMT scores. Therefore, for anyone planning to use web-mined corpora, our recommendation is to first filter out noisy sentences using heuristics and then to do ranking on the embeddings derived from the multiPLM. In contrast to the recent PDC work Steingrimsson (2023); Minh-Cong et al. (2023) that employs several deep learning model-based rigorous filtration pipelines, our technique is much simpler.
Human evaluation indicated that even after applying heuristics, some noise remains. In future, we plan to apply techniques such as classification-based approaches to remove such noise, or to apply translation post-editing to improve the sentence quality.
@@ 8.1 Limitations
We found that the LID was suboptimal for identifying Si or Ta languages. Therefore, such models would need to be optimised or better LID models would need to be used, such that they are effective in predicting the language label correctly. Furthermore, due to the lack of human annotators, we had to limit this analysis to only three language pairs. We can extend this study to more low-resource language pairs provided that we can find suitable annotators.
@@ 8.2 Ethical Concerns
We use publicly available datasets. Due to the dataset size, we did not have the resources or funding to check all the curated parallel sentences manually for offensive content. Fernando et al. (2020) provided the human-crafted dataset to be used in this research. Further, we do not disclose or share any personal details of the annotators publicly or with any other party. We have offered the annotators the standard rate and have settled the payments in full. The only details we disclose are in Appendix A.
@@ Acknowledgments
This research was funded by the Google Award for Inclusion Research (AIR) 2022 received by Surangika Ranathunga and Nisansa de Silva. We would also like to thank and acknowledge the National Languages Processing Centre (NLPC), at the University of Moratuwa for providing the GPUs to execute the experiments related to the research.
Appendix A Human Evaluation
Annotator Experience Qualification
Annotator 01 22 Diploma in Translation And Interpretation
Annotator 02 9 BSc.(Hons) in Information Technology
Annotator 03 5 MBBS
Annotator 04 4 BA (Hons) in Translation
Annotator 06 2.5 Diploma in Translation And Interpretation
Annotator 07 2.5 BSc Eng (Hons) Electrical & Electronics Engineering
Annotator 09 2 Bachelor of Industrial Information Technology
Resources Provided and Training: All the annotators have had prior experience with a similar task Ranathunga et al. (2024). However, for this annotation work, we provide them with the definitions of the noise categories (Table 9) along with example sentence pairs (Table 10) and the guideline in terms of a flowchart (Figure 3). First, we asked them to do a sample of 30 sentences as a training on the task, and review it. Then, the 1200 sentence pairs to be annotated were shared with each annotator via Google Sheets to be completed.
Parallel Corpus Categorization Code: Description
@@ Appendix A Human Evaluation
CCN: High-overlapping non-translatable text
@@
Table 9: Improved  Ranathunga et al. (2024)’s taxonomy to categorize parallel sentences to identify biases in multiPLMs.
Table 10: Example parallel sentences from the En-Si, En-Ta and Si-Ta, identified during human evaluation. The translation error in the language pair is highlighted in pink.
Compensation: They were paid the standard rate in Sri Lanka for each sentence pair they annotated or were offered co-authorship of this paper.
Appendix B Improved Noise Taxonomy
In this research, we extend the taxonomy of  Ranathunga et al. (2024) by introducing the noise category CCN (high-overlapping non-translatable text) to capture the type of noise that multiPLMs are biased to rank highly. Table 11 shows examples for this noise type. The definitions of all the categories in our improved taxonomy are given in Table 9. Then, in Table 10, we show examples of parallel sentences which fall under each of these categories.
Appendix C Sentence Length Threshold
To determine the optimal threshold for sentence length filtering, we filter sentences with fewer than 3, 4, and 5 words on the source (S) side, target (T) side, and both sides. Similar to other PDC experiments, we then rank the parallel sentence pairs by computing embeddings using XLM-R, LABSE and LASER. Finally, we train NMT models using the top 100k-ranked data. We conduct these experiments using the CCAligned corpus for En-Si and En-Ta language pairs. The ChrF++ results are shown in Table 12. We observe sentence length threshold of 5 yields the best results consistently. Therefore, we select this threshold in all subsequent experiments.
@@ Appendix C Sentence Length Threshold
Appendix D Selection of Languages and Datasets
Sinhala
is an Indo-Aryan language spoken primarily in Sri Lanka by the Sinhalese majority. It exhibits complex morphological structures, including rich inflectional and derivational processes, but is classified as a low-resource language due to the scarcity of linguistic resources and tools de Silva (2025); Ranathunga and de Silva (2022).
Tamil,
a Dravidian language with a rich literary history, is spoken by Tamil communities in Sri Lanka, India, and the global diaspora. Unlike Sinhala, Tamil benefits from a relatively larger digital presence, but it still faces challenges in NLP applications due to morphological complexity, agglutinative grammar, and resource limitations in certain domains Wijeratne et al. (2019).
We obtain the publicly released CCMatrix and CCAligned datasets from the OPUS collection Tiedemann and Thottingal (2020) [^12] https://opus.nlpl.eu/. Both these datasets support the language pairs, En-Si, En-Ta and Si-Ta, which we consider in this research.
Appendix E Selection of multiPLMs
LASER3 Heffernan et al. (2022) (L=12, H=1024, A=4, P=250M) [^14] No. of Layers, Hidden Layer Dimensions, No of Attention Heads, and Total number of parameters are defined by L, H, A, and P respectively. is a multiPLM favourable for bitext mining and cross-lingual tasks. It improves over previous LASER2 versions by supporting more languages and enhancing alignment quality, but still faces challenges in low-resource settings.
XLM-R Conneau et al. (2020) (L=12, H=768, A=6, P=278M) is a transformer-based multiPLM trained on massive amounts of text using masked language modelling. It achieves strong cross-lingual performance but struggles with low-resource languages due to limited training data.
LaBSE Feng et al. (2022) (L=12, H=768, A=12, P=471M) is a BERT-based model optimised for multilingual sentence embeddings and bitext retrieval. It provides high-quality cross-lingual representations and is favourable for cross-lingual tasks.
Appendix F NMT Experiments
The experiments are conducted on a NVIDIA Quadro RTX6000 GPU with 24GB VRAM. The hyperparameters used during training, along with the training parameters, are shown in Table 13. We conduct training on the NMT systems for 100 epochs with early stopping criteria and report the results using ChrF++. ChrF++ was chosen over the conventional multi-BLEU Papineni et al. (2002) and sacreBLEU (Post, 2018) because character-level evaluation is more suitable for the considered languages, Sinhala and Tamil, which are morphologically rich in nature.
@@ Appendix F NMT Experiments
(a) En-Si
(b) En-Ta
(c) Si-Ta
Appendix G Results Analysis
We show the individual and the combined heuristic which produced the best NMT gains with respect to each multiPLM, dataset and the language pair in the Table 14. The Figure 4 shows the percentage of dedup+n-gram experiments exceeding the highest dedup, with respect to the language-pair. Finally, we show the final dataset sizes after applying the heuristic along with the percentage reduction in Table 15.